OpenStudy (anonymous):

What comes to mind when you see this? \[ f^n(x)=(-1)^n \frac { n!}{x^{n+1}}\]

OpenStudy (anonymous):

alternating series

OpenStudy (anonymous):

this is a function sequence

OpenStudy (anonymous):

pizza....???

OpenStudy (anonymous):

Gyros?

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

ok so how does it relate to Taylor series...someone gave me a mathematical explanation, but in words, how does it relate?

OpenStudy (anonymous):

\[f^n(x)=(-1)^n \frac { n!}{x^{n+1}}\] \[\frac {f^n(a)}{n!}(x-a)^n =(-1)^n \frac {(x-a)^n}{a^{n+1}}\] \[\sum_{n=0}^\infty (-1)^n \frac {(x-a)^n}{a^{n+1}}\] I can see the math , but I'm geniunly trying to understand the concepts

OpenStudy (anonymous):